1. The electromotive force for an electric circuit with a simplified generator

is given by 𝑓(𝑑) = 40 sin 150 πœ‹π‘‘ where 𝑓(𝑑) Volts is present at 𝑑 seconds.

(a) Identify the amplitude, period, and frequency of the function.

(b) Find the maximum and minimum voltages


1
Expert's answer
2021-02-03T04:59:01-0500
  • The amplitude of a "\sin x" function is "1" and so as "f(t)=40\sin(150\pi t)", the amplitude is 40 (the expression in the argument of a sin function does not affect it's amplitude). The period of a "\sin x" function is "2 \pi", we will use it to find the period of "f(t)" : "\sin(150\pi t+2\pi)=\sin(150\pi t)", now by factoring the expressions in the parentheses we find "f(t+\frac{1}{75})=f(t)" and so the period is "T=\frac{1}{75}". The frequency is "\nu = \frac{1}{T}=75".
  • The maximum and minimum of a "\sin x" function are "-1, 1" and thus the minimum voltages are "40V, -40V" (and they are attained at, for example, "t_+=\frac{1}{300}, t_-=\frac{1}{100}"). The minimum voltage in the sense of an absolute value is "0V" (and it is attained at, for example "t_0=0").